A SEARCH FOR INTERSTELLAR H_{2}DO^{+}

\underline{CHRISTOPHER P. MORONG}, Department of Chemistry, University of Chicago; SHUJI SAITO, TAKASHI FURUYA, Research Center for Development of Far-Infrared Region, Fukui University, Fukui, Japan; HIROKO SHINNAGA, Caltech Submillimeter Observatory; TAKESHI OKA, Department of Astronomy and Astrophysics, and Department of Chemistry, University of Chicago.

The recent laboratory discoveries of submillimeter wave spectra of deuterated hydronium ions, H_{2}DO^{+}^{a} and HD_{2}O^{+},^{b} have opened up the possibility to search for these fundamental deuterated molecular ions in dense clouds where deuterium fractionation occur efficiently. We have used the Caltech Submillimeter Observatory to search for the 0_{00}^{+} → 1_{10}^{+} transition at 673.257007 GHz in IRAS16293A and Oph D. So far the freezing of CO and N_{2} has been the main issue of the depletion and deuterium fractionation, but this work adds new information on depletion and deuteration chemistry of O and O_{2}. Our observational results and analyses will be presented.

aT. Furuya, S. Saito, & M. Araki, J. Chem. Phys. 127, 244314 (2007)

bT. Araki & S. Saito, J. Chem. Phys. 128, 034311 (2008)